### v1

# arXiv:2603.01209v1[cs.AI] 01 Mar 2026

Abstract
Tool-augmented LLM agents increasingly solve tasks by interleaving natural-language reasoning with executable Python actions. Many agent frameworks equip models with a persistent interpreter, but training traces often leave this execution contract implicit. We ask whether interpreter persistence is merely a runtime scaffold, or a training-time semantic that shapes how agents learn to use the interpreter.
We answer with a controlled $2\times 2$ study, conducted on a single task family and base model, crossing training semantics (fine-tuning on persistent vs. stateless-execution traces) and runtime semantics (evaluating in a persistent vs. stateless interpreter). We introduce Opaque Knapsack, a deliberately non-collapsible, partially observable knapsack variant with budgeted tool access. For each instance, we generate matched trajectories differing only in the persistence/reset execution contract, fine-tune identical models on each variant, and evaluate all four train-runtime combinations.
We find that misaligning training and runtime semantics produces two characteristic failure modes: a model trained on persistent traces but deployed in a stateless interpreter triggers missing-variable errors in ${\sim}80\%$ of episodes, entering cascading recovery loops that consume its token budget without making progress. The reverse—a stateless-trained model deployed in a persistent runtime—pays an amnesia tax: it redundantly re-derives state that the interpreter could have retained, using roughly $3.5\times$ more tokens than the aligned persistent condition.
Notably, these efficiency and stability costs are not accompanied by a statistically significant reduction in solution quality—execution semantics primarily affect how agents reach solutions, not whether they do. This suggests that aligning fine-tuning traces with the deployment runtime reduces inference footprint without sacrificing task performance, and that interpreter persistence should be treated as a first-class semantic of agent datasets rather than a hidden implementation detail at runtime.
Keywords Tool-Augmented Language Models $\cdot$ Synthetic Training Data $\cdot$ Training–Inference Alignment
1 Introduction
Tool-augmented language models are increasingly deployed as agents that solve tasks by interleaving natural-language deliberation with executable actions—often Python code executed in an interpreter Wang et al. (2024b); Roucher et al. (2025); Gao et al. (2023)—and then reacting to environment feedback Yao et al. (2023). In these systems, the interpreter is not just a calculator: it is a workspace where the agent can accumulate variables, data structures, and intermediate results across turns.
This execution contract varies across agent frameworks. CodeAct-style agents typically run inside a persistent Python session Wang et al. (2024b), while ReAct-style agents externalize state into the context window Yao et al. (2023), and other systems maintain dedicated memory stores Packer et al. (2024). Yet training pipelines usually treat agent traces as token sequences and leave execution semantics implicit. This mismatch motivates our research question:
Is interpreter persistence merely an inference-time scaffold, or a training-time semantic that shapes how agents learn to use tools?
The distinction matters in practice: Agent builders routinely fine-tune models on tool-use logs produced by a particular runtime and then deploy the resulting model in a (sometimes different) runtime. If execution semantics are learnable, then matching training traces to deployment semantics can determine whether a CodeAct agent exploits persistent state and how efficiently it does so under latency and cost constraints.
We answer this question with a controlled $2\times 2$ study that disentangles training semantics from runtime scaffolding, conducted on a single task family and base model. We introduce Opaque Knapsack, a partially observable knapsack variant designed to be non-collapsible for executable-action agents: Item attributes and feasibility constraints are hidden behind budgeted tool calls, and per-step interaction limits force iterative information gathering and state revision. For each task instance, we generate paired interleaved reasoning-action trajectories matched by task instance, prompt, tool interface, and supervision, differing only in the persistence/reset execution contract: whether the interpreter state persists across turns or is reset after each action. We fine-tune identical Qwen3-8B models on each trace type and evaluate them under both runtime semantics.
The resulting cross-evaluation shows that persistence is not a strictly zero-shot capability: it must be learned. In terms of mean normalized optimality, we do not observe statistically significant differences between the fine-tuned models across training/runtime semantics at $n=100$ tasks per split (Appendix E); we therefore treat score differences as directional. However, execution semantics substantially change how agents use the interpreter: When the deployment runtime is persistent, persistent-trained agents reuse executable state and complete episodes with far fewer tokens; stateless-trained agents redundantly externalize state into text across all deployments—an “amnesia tax” that persists even when a persistent runtime is available; and when a persistent-trained model is deployed under a stateless runtime, it exhibits characteristic state-mismatch failures (missing-variable exceptions) and instability. Overall, execution semantics matter at train time—not only as a post-hoc execution harness—because they shape the learned state-management strategy and its brittleness under deployment.
Contributions:
  • Execution semantics as a training variable. We identify interpreter persistence as part of the agent trace contract, studying execution semantics as learnable components of the training signal with a controlled $2\times 2$ train/runtime design.
  • A non-collapsible benchmark with paired trajectories. We contribute Opaque Knapsack and a paired trace generation pipeline. This benchmark toggles only the persistence contract while holding tasks, tools, and supervision fixed, thus forcing multi-turn control flow and iterative state revision in tool-augmented agents.
  • Evidence that persistence is learned. Using a $2\times 2$ train/runtime cross-evaluation, we show that exposure to persistence during fine-tuning shapes state-management behavior. These results suggest that interpreter persistence should be treated as a first-class semantic of agent datasets and training protocols, not a hidden implementation detail of the runtime. Relatedly, aligning the execution semantics of training traces with the deployment runtime lowers inference cost without compromising performance.
2 Related Work
Tool-Augmented Agents and Agent–Computer Interfaces.
A growing body of work studies agents that interleave natural-language reasoning with executable actions. Program-aided and executable-action frameworks such as PAL (Gao et al., 2023), PoT (Chen et al., 2023b), ToRA (Gou et al., 2024), and CodeAct (Wang et al., 2024b; Roucher et al., 2025) demonstrate that offloading computation to an external interpreter can improve reliability on multi-step tasks. Interleaved reasoning–action paradigms such as ReAct (Yao et al., 2023) further formalize agents as systems that alternate between deliberation and interaction. Recently, the Agent–Computer Interface (ACI) perspective has emphasized that the interface between an agent and its execution environment—including tool affordances, feedback formats, and an interactive runtime—is itself a key design surface for agent performance. For example, SWE-agent (Yang et al., 2024) shows that purpose-built ACIs for software engineering tasks substantially improve agent effectiveness. However, ACI work largely treats the execution environment as an inference-time systems choice. In contrast, we isolate a specific execution semantic—interpreter state persistence—and study it as part of the training-time trace contract, measuring how training/runtime alignment shapes learned state-management behavior.
Stateful and Non-Collapsible Evaluation Environments.
In many tool-use settings, if task-relevant information is fully exposed (or tools behave as purely stateless functions), an agent can sometimes solve instances with a single long script, reducing the need for multi-turn interaction and cross-step state tracking. Recent benchmarks have begun to emphasize stateful tool execution and longer-horizon dependencies. ToolSandbox (Lu et al., 2025), for example, contrasts stateless service calls with stateful tool execution, requiring agents to handle implicit dependencies that unfold across turns. Similarly, InterCode (Yang et al., 2023) evaluates agents in interactive code execution environments, though execution semantics are typically fixed rather than manipulated. Our Opaque Knapsack environment builds on this motivation: by combining hidden constraints, budgeted tool access, and strict partial observability, it enforces iterative information acquisition and plan revision over multiple turns. This design makes the task sensitive to where state lives—as executable bindings in a persistent interpreter versus as text reconstructed from the context window—without assuming persistence is strictly required for solvability.
Execution Semantics as a Blind Spot in Agent Training Data.
Synthetic interaction traces are a primary source of supervision for agentic systems, and numerous pipelines focus on trajectory quality, verification, and diversity. Process supervision shows that supervising intermediate reasoning can improve robustness (Lightman et al., 2023), motivating multi-turn trajectory synthesis via simulated environments, verification loops, failure-driven sampling, and self-improvement (Hao et al., 2026; Prabhakar et al., 2025; Zelikman et al., 2022). Agent-specific fine-tuning frameworks such as FireAct (Chen et al., 2023a) and AgentTuning (Zeng et al., 2023) demonstrate that SFT on such trajectories can instill general tool-use behaviors. Concurrently, schema efforts like the Agent Data Protocol (ADP) (Song et al., 2025) aim to standardize actions and observations, and execution-centric training approaches leverage program traces for verification and grounded supervision (Jin et al., 2025; Jung et al., 2025). Across these lines of work, the dominant abstractions specify what tools are called and what observations are emitted, but typically leave the underlying execution semantics implicit—in particular, the lifetime of interpreter bindings and how executable state evolves across steps. Our work targets this omission directly by treating interpreter persistence as a controlled semantic of the data-generation and training protocol.
Training–Inference Behavioral Alignment.
A recurring theme in agent learning is that capabilities needed at deployment must appear in the training distribution. Toolformer (Schick et al., 2023) shows that effective tool invocation emerges when tool use is embedded in training, and recent work argues that the allocation and timing of reasoning should align between training and inference (Nguyen et al., 2026; Akter et al., 2025). We extend this alignment principle to execution semantics: our results suggest that if persistent state is available at deployment, exposing the same persistence contract during fine-tuning can shape whether (and how) agents learn to delegate state to the interpreter. Conversely, mismatching training trace semantics and deployment runtimes can induce compounding behavioral distribution shift, a known vulnerability in imitation learning (Ross et al., 2011).
Explicit State, Context Growth, and Efficiency.
Long-horizon agents often externalize intermediate state into text (e.g., via scratchpads), which increases context length as interaction histories grow. While some approaches investigate implicit (latent) reasoning to reduce latency (Deng et al., 2023), our study highlights a complementary efficiency axis for tool-augmented agents: when persistence is available and learned, executable bindings can serve as a compact external memory, reducing redundant re-derivation and re-expression of state in the context window (the “amnesia tax”).
3 Methodology
Figure 1: Interpreter persistence as an execution semantic for tool-augmented agents. (Top) Opaque Knapsack hides item attributes and feasibility constraints behind a budgeted tool API, forcing multi-turn inspection and plan revision rather than a single one-shot script. (Bottom) Example rollout for the same task instance under two deployment runtimes. With a persistent interpreter (left), variables defined by earlier actions remain live and the agent can accumulate and reuse executable state across turns (e.g., items, inspect_data). With a stateless interpreter (right), globals are cleared after each step while the text history remains; a policy trained under persistence may still reference prior bindings, producing missing-variable (NameError) failures and costly reconstruction loops. We quantify how these behaviors depend on train–runtime alignment in a controlled $2{\times}2$ study (persistent vs. stateless traces $\times$ persistent vs. stateless runtime).
We study whether persistent executable state functions as a learnable inductive bias as opposed to an inference-time scaffold. Accordingly, our methodology focuses on the structure and the execution semantics of agent training data.
Defining Execution Semantics.
We use execution semantics to denote the operational contract implemented by the agent runtime: a transition function
$\mathcal{E}:(a_{t},s_{t})\mapsto(o_{t},s_{t+1}),$
mapping an emitted Python action $a_{t}$ and current runtime state $s_{t}$ to the next observation $o_{t}$ and next runtime state $s_{t+1}$. This contract specifies (i) the lifetime of interpreter bindings across steps, (ii) how tool calls are executed and exceptions are surfaced, and (iii) how outputs and observations are serialized into the context. We hold the task distribution, trace format, tools, and supervision fixed — as well as (ii)–(iii) above — and vary only (i): whether executable state persists across agent steps or is reset after each action.
To isolate and study this single dimension of execution semantics, we first introduce our non-collapsible task environment (Section 3.1), describe the generation of interleaved reasoning trajectories (Section 3.2), and formally define the persistent and stateless regimes (Section 3.3).
3.1 Task: Opaque Knapsack
For our controlled study, we design Opaque Knapsack, a procedurally generated, partially observable optimization task based on the classical 0/1 knapsack problem (Kellerer et al., 2004), designed to require multi-step interaction and state maintenance (Figure 1). Each instance presents a set of items with hidden attributes (weight, value, class) and a hidden allowed-class constraint, making the task partially observable in the standard POMDP sense (Kaelbling et al., 1998): the agent receives observations that are an incomplete function of the underlying task state. Item attributes are accessible only via a budgeted inspection tool; class validity is not directly revealed and must be inferred from interaction feedback (e.g., take_item successes/errors). The agent must select a value-maximizing subset within a capacity limit, and must inspect an item before taking it.
By non-collapsible we mean the task cannot be solved by a single open-loop script. Three properties enforce this: (i) attributes are tool-gated, forcing incremental information acquisition (Wang et al., 2024c; Yao et al., 2024); (ii) feasibility depends on hidden class constraints and may invalidate previously inspected high-value candidates, forcing plan revision; and (iii) inspection is budget-limited, preventing exhaustive enumeration (Liu et al., 2025, 2026; Ding et al., 2026). Together, these constraints require genuine interleaving of reasoning, execution, and observation across turns (Jiang et al., 2026), making the task particularly sensitive to whether executable state persists across actions. Agents interact through a three-tool task API (list_items, inspect, take_item); full task formalism and API details are in Appendix A.
3.2 Interleaved Reasoning Trajectories
We fine-tune models on interleaved reasoning–execution trajectories generated by a CodeAct-style teacher Wang et al. (2024b); Roucher et al. (2025). Each trajectory strictly alternates between a brief natural-language reflection and a single executable Python action, followed by environment observations Yao et al. (2023). Actions may invoke tools, perform control flow, and update interpreter variables, with all execution outputs appended to the context for subsequent steps. While trajectories are matched by task instance, tools, and base supervision, they differ in the specific runtime instructions and format-only few-shot demonstrations required to faithfully reflect the persistent or stateless execution contract.
3.3 Persistent vs. Stateless Execution
For each training instance, we generate two execution variants that differ only in how interpreter state is handled across steps. In the persistent regime, variables and data structures defined in one action remain available in subsequent actions. In the reset regime, interpreter state is cleared after every step, requiring the model to re-derive any required executable state.
Importantly, stateless execution does not remove observations from the context window; it removes only the ability of the interpreter to carry mutable state across steps. Thus, both regimes expose identical observations and supervision and differ solely in execution semantics.
The full CodeAct-style loop, runtime-state header, and enforcement contract are described in Appendix B.
4 Empirical Study: Opaque Knapsack Experiments
We provide an empirical validation designed to isolate the role of state persistence in multi-step interleaved reasoning under partial observability.
4.1 Experimental Setup
This section details the end-to-end experimental pipeline designed to isolate and evaluate the role of execution semantics. We outline the procedural generation of the task splits and teacher trajectories, followed by the specific hyperparameter configurations for model fine-tuning and the infrastructure used for downstream inference.
4.1.1 Datasets and Trace Generation
Tasks and splits.
We evaluate on the Opaque Knapsack task generated procedurally with two difficulty regimes: Easy instances are used for training and in-domain evaluation, while Hard instances are held out and used exclusively for evaluation. All instances share the same tool interface and evaluation protocol. Difficulty is scaled solely by instance size and constraint tightness. For training, we sample 1,000 Easy instances, and for evaluation we sample 100 tasks for each of the Easy and Hard configurations. Full generation ranges, class sampling, rejection constraints, and solver-based filtering are detailed in Appendix A.
Trace generation.
Training traces are generated using a persistent state runtime, instantiated as a CodeAct-style teacher agent that alternates brief natural-language reflection with a single executable Python action per turn. The teacher interacts with each task environment exclusively through declared tools, and all code blocks are executed in a sandboxed interpreter, with resulting observations appended to the context. We use a single teacher model (Gemini 3 Flash Google DeepMind (2025)) and generate trajectories under a fixed turn and tool-call budget. To isolate the role of executable state, we generate two trace variants with identical tasks and prompts: a persistent variant where interpreter globals persist across turns, and a stateless variant where interpreter state is cleared after each step. All other generation settings are held fixed. Exact system prompts, runtime headers, and few-shot demonstrations are documented verbatim in Appendix B.
Few-shot demonstrations and contract consistency.
Each runtime regime includes a small number of format-only demonstrations whose sole purpose is to make the execution contract concrete (persistent reuse vs. mandatory redefinition / explicit printing of state). These demonstrations are intentionally task-agnostic (they do not contain knapsack-specific information) and are treated as part of the runtime interface: using persistence-style examples in a stateless interpreter (or vice versa) would create an internally inconsistent protocol. At evaluation time the prompt—including the demonstrations and runtime instruction—is held fixed within each runtime. Thus, differences we report between persistent-trained and stateless-trained models under the same runtime (e.g., token footprint under a persistent interpreter) cannot be attributed to different demonstrations at inference. Moreover, in the mismatch condition the persistent-trained model is explicitly shown stateless-style demonstrations yet still produces missing-binding errors characteristic of reliance on cross-turn executable state, indicating that the dependence is learned rather than merely prompted.
Dataset characteristics.
Although the persistent and stateless execution regimes are generated from identical task instances, prompts, and tool interfaces, they induce distinct interaction distributions. Our training sets are matched by episodes: each regime contains the same number of trajectories, generated on the same number of sampled task instances. Table 1 summarizes key properties of the resulting trace datasets. Teacher behavior is broadly comparable across regimes, with similar inspection patterns and capacity utilization, and with small differences in success rate and normalized optimality.
[…]
Metric Persistent Execution Stateless Execution
Teacher Success Rate (%) 17.8 22.8
Avg. Teacher Optimality 0.671 0.745
Avg. Capacity Utilization (%) 62.6 71.2
Avg. Items Inspected 27.4 27.5
Avg. Steps per Episode 4.31 6.55
Avg. Tool Calls per Episode 8.39 11.54
Avg. Total Tokens per Episode 18,337 55,516
Teacher Success Rate is the fraction of episodes in which the teacher agent achieved exactly the instance-optimal value (normalized optimality =1.0=1.0), equivalent to the Solved column in Table 2.
Table 2: Opaque Knapsack evaluation across training and runtime execution semantics. We report task performance (normalized optimality and exact solves) alongside computational footprint (steps, total tokens, and wall-clock time) for Easy and Hard splits. Base model results are in gray. Margins ($\pm$) denote 95% percentile bootstrap Efron and Tibshirani (1993) confidence intervals over the $n=100$ evaluation task instances (5,000 resamples). Score / 1k Tokens measures efficiency as the score achieved per 1,000 total tokens. The results demonstrate that interpreter persistence is a highly effective, learnable semantic: the persistent-trained model operating in a persistent runtime achieves comparable optimality to the stateless baseline while consuming roughly one-third as many tokens. Conversely, stateless-trained agents pay an “amnesia tax” regardless of runtime, redundantly externalizing state even when the interpreter could retain it.
[…]
Tokens: cumulative total tokens (prompt ++ completion) across all turns; grows with context as conversation history accumulates.
In contrast, the interaction footprint differs substantially. Stateless execution produces longer trajectories with more steps per episode, more tool calls, and more total tokens per episode on average. As a result, matching by episodes implies that the stateless dataset contains substantially more total tokens and tool interactions than the persistent dataset. We report downstream model performance and cross-evaluation results separately in Sections 4.3 and full statistical tests in Appendix E.
Token volume and trace quality.
Because we match training sets by episodes (same task instances; same number of trajectories), the stateless-execution traces are substantially longer in raw token count (Table 1). This length difference does not imply proportionally more distinct supervision: the number of inspected items is essentially identical across regimes (27.4 vs. 27.5), and the additional tokens in the stateless regime are largely induced by the need to repeatedly reconstruct and re-express intermediate state that would otherwise persist in the interpreter, consistent with the state-shift measured in Figure  3. Moreover, the stateless teacher is not weaker on outcome metrics-if anything it is slightly stronger (higher success rate and normalized optimality; Table 1)-so the persistent condition is not advantaged by an obvious teacher-quality explanation. We therefore interpret raw token count as a poor proxy for learnability in this setting, and treat small differences in normalized optimality as directional; the most stable effects we observe are instead semantic and conditional on train–runtime alignment.
4.1.2 Model Fine-Tuning and Inference
Training data preparation. Raw trace logs are converted into chat-style training examples via a deterministic four-stage pipeline: (i) structural and outcome validation to filter degenerate trajectories Wang et al. (2025, 2024a); (ii) message extraction into the standard system, user, and assistant chat format Pan et al. (2026); (iii) context-aware truncation to the model’s sequence limit, preserving causal coherence Drouin et al. (2024) by retaining the task header and final action while backfilling recent turns; and (iv) filtering traces that cannot be coherently truncated. Persistent and stateless traces are processed under identical policies; full data preparation, fine-tuning on the Qwen3-8B model, and inference details are provided in Appendix C.
4.2 Evaluation Protocol and Metrics
We evaluate a $2\times 2$ design that crosses training-time execution semantics with runtime execution semantics. Specifically, we fine-tune two LoRA adapters on matched sets of $1,000$ Opaque Knapsack episodes of Easy difficulty (as defined in Appendix A): a persistent-trained adapter trained on traces generated under persistent execution, and a stateless-trained adapter trained on traces generated under stateless execution. At evaluation time, each adapter is run under both a persistent and a stateless interpreter runtime, yielding four experimental conditions.
All conditions are evaluated on the same fixed set of 100 Easy (in-domain) and 100 Hard (scaled-difficulty) Opaque Knapsack instances. Task solution performance is measured by normalized optimality (achieved value divided by the instance optimal value), and we additionally report the number of exact solves (instances achieving the optimum). For each episode, we also record interaction footprint (steps, total tokens, and wall-clock time). Statistical significance is assessed using paired two-sided Wilcoxon signed-rank tests Wilcoxon (1945) over per-instance normalized optimality scores (Appendix E). We also report Score / 1k Tokens—normalized optimality divided by total tokens generated per episode (in thousands)—as a joint measure of solution quality per unit inference cost; higher values indicate greater efficiency.
4.3 Experimental Results
Table 2 summarizes performance on 100 Easy (in-domain) and 100 Hard (scaled-difficulty) Opaque Knapsack instances.
Efficiency Tradeoff and Interaction Footprint.
Figure 2: Cost–performance tradeoff under persistent vs. stateless runtime. The y-axis reports mean normalized optimality (percentage of the ground-truth optimal value). The x-axis reports total tokens generated per episode (log scale). Removing runtime persistence substantially increases token cost with modest impact on normalized optimality.
Although Easy-task performance is similar across regimes, the computational footprint differs substantially. On the Hard split, the end-to-end stateless configuration requires approximately $3.5\times$ more tokens than the end-to-end persistent configuration (67,898 vs. 18,612 tokens). We term this overhead the amnesia tax—the token cost of iteratively re-deriving and externalizing state when the interpreter cannot retain it.
[…]
Execution Footprint Analysis: State Mismatch and Execution Instability.
We next analyze trace-level behavior to understand how persistent training shapes execution dynamics.
When the persistent-trained model is evaluated under stateless runtime, state-missing exceptions emerge at high frequency. On both Easy and Hard splits, approximately 80% of traces contain at least one unresolved reference error, with an average of 1.28 such exceptions per Hard episode. These errors arise when the model attempts to reference runtime variables defined in previous turns—an assumption valid under persistent execution but invalid under stateless semantics.
No unresolved reference errors are observed in other train/runtime configurations. This asymmetry provides direct evidence that persistent training induces genuine reliance on cross-turn executable state. Notably, this vulnerability occurs despite the runtime explicitly exposing active_globals and last_step_globals in the observation header (Appendix B). Because this variable manifest makes the missing state fully observable to the agent, one might expect the mismatched model to detect the anomaly and adapt. The fact that it still collapses into cascading execution failures makes the depth of this learned persistence prior all the more surprising.
The exact trace pairing, event parsing, and error-density thresholds are described in Appendix D.
Failure Mode Analysis: Execution Instability vs. Clean Suboptimality.
To understand the effect of deploying a persistent-trained model under a stateless runtime, we analyze terminations on the Hard split.
Among runs where the task was not solved successfully:
  • 11.6% terminate due to turn or context exhaustion,
  • 12.8% exhibit protocol or sequencing violations,
  • 57% display cascading execution failures,
  • 18.6% fail silently without explicit errors.
We define a cascading execution failure operationally as episodes exhibiting either a high error-to-step ratio ($50\%$) or repeated error concentration in the final interaction window. These trajectories typically involve repeated recovery attempts following state-mismatch exceptions, leading to cascading reconstruction loops that degrade effective reasoning progress.
The dominant failure mode is instability introduced by learned assumptions about persistent state, not Knapsack rule violation. When the runtime removes persistent state, the model is pushed into scaled-difficulty recovery states that the persistent teacher never demonstrated. This vulnerability to compounding errors when the runtime distribution shifts away from the training distribution is a known limitation of standard imitation learning and behavioral cloning policies (Ross et al., 2011). Consequently, the model frequently attempts to recover but enters high-error-density loops, which become the dominant source of suboptimality in this condition. Failure classification categories and decision order are defined formally in Appendix D.3.
Table 3: Trace-level diagnostics on Hard difficulty tasks ($n=100$). Among fine-tuned adapters, Unresolved reference error incidence is an exclusive signature of train–runtime mismatch: It appears only when the persistent-trained model is executed under a stateless runtime. Unresolved reference error statistics are computed across all episodes. The failure-mode breakdown conditions on normally-terminated non-optimal episodes under the diagnostic taxonomy (Appendix D.3).
Train $\rightarrow$ Runtime % Unresolved Ref. Errors Unresolved Ref. Errors / Episode Normally Terminated (Non-Optimal) Constraint / Protocol Violation Execution Instability Silent Suboptimality
Persistent $\rightarrow$ Persistent 0% 0.00 91 13 34 44
Persistent $\rightarrow$ Stateless 80% 1.28 76 11 49 16
Stateless $\rightarrow$ Persistent 0% 0.00 86 7 1 78
Stateless $\rightarrow$ Stateless 0% 0.00 74 10 0 64
5 Discussion
Our results support a practical claim: Execution semantics observed during training shape how a tool-augmented agent learns to use the interpreter at deployment. In a controlled $2\times 2$ cross-evaluation (training semantics $\times$ runtime semantics; Table 2), models adapt their state management strategy to the persistence contract embedded in their fine-tuning traces. When the deployment runtime exposes a persistent interpreter, aligning fine-tuning traces with that same contract produces agents that more effectively reuse executable state (Figure 2), with measurable gains in interaction efficiency and directionally higher solution quality.
Training–runtime alignment is a lever for agent builders.
Agent frameworks differ in whether state is carried by the interpreter (CodeAct-style persistence) or pushed into the context window (stateless/textual state). Our findings suggest that this is not just an implementation detail of the harness: it is a learnable behavioral prior. Fine-tuning on traces that reflect the intended runtime contract can directly shape an agent toward desirable behaviors such as stable variable naming, incremental decomposition across turns, and in-place updates of shared data structures. In settings where agents are scrutinized for both performance and cost, this alignment becomes actionable: On the harder split, the persistent-trained model under a persistent runtime achieves comparable mean normalized optimality (75.4% vs. 72.5%, a difference that is not statistically significant at $n=100$; Table 8) while using far fewer tokens per episode than a stateless-trained model in the same runtime (Table 2).
Persistence changes where state lives.
Figures 2 and 3 highlight a core trade-off. Stateless training shifts the locus of state from the interpreter into the context window, inducing what we term the “amnesia tax”: the token cost of iteratively re-deriving and externalizing state the runtime cannot retain. Critically, this tax is a learned behavioral prior, not a runtime response. Even when deployed in a persistent interpreter, the stateless-trained adapter shows State Utilization =0=0 and Imports/Step =1.0=1.0 (Appendix F)—the model re-imports and re-derives on every step despite the runtime being fully capable of retaining those bindings.
Conversely, persistent execution makes it advantageous to keep task-relevant state as executable bindings with non-zero lifespan in the interpreter, reducing redundancy and enabling shorter, more stable interaction traces. Importantly, the stateless regime is not data-limited (as shown in Table 1): It produces longer, more tool-heavy trajectories and substantially more tokens during trace generation, yet yields a substantially larger interaction footprint and directionally lower solution quality downstream. This indicates that token volume is an unreliable proxy for learnability in agent datasets; the execution contract can make traces qualitatively easier or harder to learn from even when tasks, tools, and supervision are held fixed. This echoes findings that SFT teaches behavioral style and format as much as task capability (Zhou et al., 2023) — here, the execution contract is the stylistic prior being absorbed.
Mismatch failures reveal genuine reliance on executable state.
Misalignment between what the model was trained to assume and what the runtime actually provides induces characteristic failures. When a persistent-trained model is evaluated under a stateless runtime, it produces missing-binding exceptions in roughly $80\%$ of episodes and frequently enters error-dense recovery loops (Section 4.3). These failures are asymmetric: They do not appear in the other train/runtime configurations, indicating that persistent-trained policies are not merely “helped” by statefulness but depend on it as part of their learned execution strategy. From an engineering perspective, this suggests a concrete risk: swapping runtimes (or silently changing interpreter persistence) can degrade both correctness and stability without any change to the model weights.
Implications for dataset and system design.
Two implications follow: First, execution semantics (e.g., persistence vs. statelessness as well as how state is exposed or summarized) should be treated as first-class attributes of agent datasets and training protocols, alongside tool schemata and supervision signals. Second, benchmarks for executable-action agents should explicitly prevent collapse into one-shot scripts; Opaque Knapsack provides one such blueprint by combining tool-gated hidden attributes, hard interaction budgets, and non-monotonic constraints that force iterative revision.
Falsifiability.
If interpreter persistence were only an inference-time scaffold, we would expect little or no interaction between training semantics and runtime semantics in our design. Specifically, a persistent runtime should reduce token cost similarly regardless of which adapter is deployed; a persistent-trained model should transfer cleanly to a stateless interpreter when shown the stateless contract, without systematic missing-binding errors; and the base model should benefit materially from persistent execution once tool access is available. Any of these outcomes would weaken the conclusion that persistence is a learned behavioral prior that must be aligned between training traces and the deployment runtime. Instead, we observe the opposite in each case: Token efficiency gains are concentrated in the aligned persistent condition, missing-binding errors are asymmetric and specific to the persistent-trained/stateless-runtime mismatch, and the base model remains near-floor regardless of runtime statefulness—which is consistent with persistence functioning as a training-time semantic rather than a post-hoc scaffold.
Limitations and future work.
We highlight four limitations of our current study that should be addressed in follow-up work.
  • Evaluation power: While our sample size ($n=100$ tasks per split) cleanly isolates massive behavioral and efficiency shifts, it is underpowered to conclusively resolve differences in absolute solution optimality.
  • Token-budget confounds: Matching training sets by episode rather than total tokens exposed the stateless-trained model to substantially more raw text ($\sim 3.5\times$). While much of this volume reflects redundant state reconstruction, future work should explicitly ablate token-matched training budgets.
  • Scope of generalization: Our controlled design cleanly isolates the variable of interest using a single task family and base model. Validating these findings across varied models and tasks is an important next step.
  • Co-occurring protocol cues: Our runtime provided structured metadata about prior-step interpreter symbols to equalize observability. Factorizing these visibility signals from pure persistence mechanics will further clarify how agents learn to manage memory.
6 Conclusion
We have shown that interpreter persistence is not merely an inference-time scaffold but a learnable behavioral prior that must be aligned between training traces and the deployment runtime. In a controlled $\times2$ cross-evaluation on Opaque Knapsack, persistent-trained agents deployed in a persistent runtime achieve comparable solution quality at substantially lower token cost, while mismatched deployments produce either cascading execution failures or a portable “amnesia tax” that persists regardless of what the runtime offers. The practical implication for agent builders is direct: when fine-tuning on tool-use traces, the execution semantics of the runtime used to generate those traces should be treated as a first-class design decision, not a hidden implementation detail. Future work should establish how broadly these findings generalize across task families, model scales, and runtimes that occupy the space between fully persistent and fully stateless execution.
Appendix A Task Specification and Generation Details
This appendix provides additional details on the procedural generation of Opaque Knapsack instances used in the empirical study (Section 4). We provide a formal definition of the task, report its API contract, full configuration for the Easy and Hard difficulty buckets and describe the constraints enforced by the sampler to ensure non-degenerate, execution-dependent instances.
Figure 4: Opaque Knapsack interaction interface and hidden task state. The agent interacts with each instance exclusively through a fixed tool API. Item attributes (weight, value, class), class validity constraints, and feasibility state are not directly observable and must be discovered or inferred through tool-mediated interaction. This partial observability and tool-gated access prevent one-shot solutions and require iterative information gathering and state maintenance.
A.1 Formal Task Definition
An Opaque Knapsack instance is defined by a tuple $(\mathcal{I},C,\mathcal{K},\mathcal{A},B)$, where $\mathcal{I}=\{i_{1},\dots,i_{N}\}$ is a set of items identified by opaque IDs, $C\in\mathbb{R}^{+}$ is the knapsack capacity, $\mathcal{K}$ is the universe of item classes, $\mathcal{A}\subseteq\mathcal{K}$ is a hidden subset of allowed classes, and $B\in\mathbb{N}$ is an inspection budget. Each item $i\in\mathcal{I}$ has latent attributes $(w_{i},v_{i},c_{i})$ corresponding to weight, value, and class. These attributes may only be accessed by invoking an inspection tool, which reveals $(w_{i},v_{i},c_{i})$ and consumes one unit of inspection budget. The agent must select a subset $S\subseteq\mathcal{I}$ maximizing the total value:
$\max_{S\subseteq\mathcal{I}}\sum_{i\in S}v_{i}\quad\text{s.t.}\quad\sum_{i\in S}w_{i}\leq C,\;c_{i}\in\mathcal{A}\ \forall i\in S,\;|\mathcal{Q}|\leq B,\;S\subseteq\mathcal{Q},$
where $\mathcal{Q}$ is the set of inspected items.
A.2 Task Interaction Interface and API Contract
Agents interact with each Opaque Knapsack instance exclusively through the three-tool interaction interface shown in Figure 4; the API is detailed in Table 4. Tool return values and execution errors are appended verbatim to the context and made available as observations in subsequent steps.
Table 4: Opaque Knapsack interaction API and hard budgets. Each instance specifies a knapsack capacity $C$ and an inspection budget $B$ that limits the number of distinct items whose attributes may be queried. Tool return values (JSON strings) and tool exceptions are appended verbatim to the context as observations.
Call $\rightarrow$ return (observation) Budgets / constraints
list_items() $\rightarrow$ JSON string of all opaque item IDs, e.g. ’["item_...","item_..."]’ No budget cost. Reveals the candidate item ID set.
inspect(item_id: str) $\rightarrow$ JSON string with keys {"class","value","weight"}, e.g. ’{"class":"A","value":13,"weight":12}’ Consumes 1 unit of inspection budget $B$ on first inspection of a given ID. Repeated inspections are cached and free. Errors if ID is unknown or if $B$ is exceeded.
take_item(item_id: str) $\rightarrow$ None Adds the item to the knapsack. Errors if the item was already taken, belongs to a disallowed class, or would exceed capacity $C$.
finish() $\rightarrow$ None Terminates the episode (runtime-provided; not task-specific).
The inspection budget $B$ is instance-specific, computed adaptively from capacity, item distribution, and optimal solution size.
A.3 Difficulty buckets
We generate two difficulty buckets from the same underlying knapsack generator, differing only in scale and constraint tightness.
Easy bucket.
Easy instances are designed to be solvable within short horizons while still requiring explicit inspection, state tracking, and combinatorial reasoning. Each instance samples: (i) 25–40 items, (ii) item weights uniformly from [5, 20], (iii) item values uniformly from [10, 100], and (iv) item classes from a fixed set of 15 symbols. A subset of 3 classes is designated as allowed. The knapsack capacity is set to a random fraction in [0.35, 0.5] of the total weight of allowed items. Easy instances are used for both training and in-domain evaluation.
Hard bucket.
Hard instances scale the same task structure to substantially larger problem sizes and tighter feasibility constraints. Each instance samples: (i) 80–120 items, (ii) item weights uniformly from [5, 50], (iii) item values uniformly from [10, 500], and (iv) item classes from a set of up to 26 symbols. A subset of 5 classes is designated as allowed. The knapsack capacity is set to a random fraction in [0.4, 0.6] of the total weight of allowed items. Hard instances are held out from training and used exclusively to evaluate difficulty generalization.
A.4 Structural Constraints and Rejection Sampling
Instances are generated using rejection sampling to enforce minimal combinatorial structure and to rule out degenerate solutions. For each candidate instance, we compute the optimal solution using a deterministic 0/1 knapsack dynamic programming solver restricted to allowed-class items. An instance is accepted only if all of the following conditions are satisfied:
  • Minimum solution size. The optimal solution must contain at least three items, ensuring that the task requires combining multiple inspected items rather than selecting a single dominant element.
  • Anti-dominance constraint. No single item in the optimal solution may contribute more than 40% of the total optimal value. This prevents instances where the task can be solved by identifying one obvious outlier.
  • Feasibility. The sampled capacity must be at least as large as the minimum weight of any allowed item, ensuring that at least one feasible solution exists.
Candidate instances failing any of these checks are rejected and resampled.
A.5 Inspection Budget Derivation
Each instance includes a strict inspection budget that limits the number of distinct items whose properties may be queried via the inspect tool. The budget is computed adaptively to balance exploration pressure with solvability.
Let $p_{\mathrm{valid}}$ denote the fraction of items belonging to allowed classes, and let $\bar{w}$ denote the mean item weight. We estimate the expected number of valid items required to fill the knapsack as $C/\bar{w}$, where $C$ is the capacity. The inspection budget is then set to exceed this estimate with a safety margin, scaled by $1/p_{\mathrm{valid}}$ to account for class filtering.
Additionally, the budget is lower-bounded by a deterministic floor proportional to the size of the optimal solution, ensuring that inspecting all items in the optimal set is always feasible. The final budget is clamped to lie between a minimum of five inspections and the total number of items in the instance.
This design ensures that agents must make strategic inspection decisions while preventing trivial exhaustive inspection.
A.6 Reference Solutions
For each accepted instance, we store the optimal item set, total value, and total weight computed by the reference solver. These reference solutions are used exclusively for evaluation and scoring; they are never revealed to the agent during trace generation or inference.
Appendix B Agent Implementation
To generate valid trajectories interleaving reasoning and executable actions, we implemented a custom CodeAct-style agent loop. This appendix documents the execution contract, runtime alignment across regimes, and the exact system prompts used in the experiments.
B.1 Execution Flow and Turn Structure
Each episode follows a strict Reflection–Action–Observation loop:
  1. 1.
    Reflection (Natural Language). The model first emits a brief explanation of the next step.
  2. 2.
    Action (Executable Python). Exactly one fenced Python code block is allowed per turn.
  3. 3.
    Observation (Environment Output). The code is executed in a sandboxed interpreter. The resulting output, return value, and any exception are serialized into a structured JSON observation and appended to the context.
Strict Turn-Taking and Format Enforcement.
The agent enforces a “one code block per turn” protocol by extracting all fenced code blocks from the model response. If at least one code block is present, the agent executes only the first block and discards any additional blocks. When multiple blocks are detected, the agent records this fact in the subsequent observation payload (as a system_note) to make the deviation explicit. If no code block is found, the agent does not execute anything and instead appends a structured error observation instructing the model to retry with a single fenced Python block.
B.2 Structured Runtime-State Header (Aligned Across Regimes)
To equalize state observability across execution regimes, every turn appends a structured JSON object of the form detailed in Listing 1:
Listing 1: The structured JSON runtime-state header appended to every turn to equalize state observability.
"observation": {
"success": bool,
"result": ...,
"output": "...",
"error": ...
"runtime_state": {
"runtime": "persistent" | "reset",
"active_globals": [...],
"last_step_globals": [...]
The field last_step_globals lists the variable names that existed after executing the previous code block in both regimes. The field active_globals lists the variable names that are currently executable in the interpreter.
  • In the persistent regime:
    $\texttt{active\_globals}=\texttt{last\_step\_globals}.$
  • In the stateless regime:
    $\texttt{active\_globals}=\varnothing,$
    while last_step_globals still reports the previous names.
No variable values are exposed to the model—only symbol names. Thus, both regimes receive identical structural metadata about the prior step. The only semantic difference is whether previously defined bindings remain executable in the interpreter.
The interpreter enforces persistence or stateless behavior independently of the prompt text. Even if the model ignores the runtime instruction, the environment either preserves or clears bindings accordingly. Therefore, the manipulated variable is the execution semantics of the runtime, not additional supervision.
B.3 Semantic Termination
Episodes terminate only when the model explicitly invokes the finish() tool inside a Python block. This ensures that each trajectory represents a complete reasoning chain where the agent determines when the task is solved.
If the maximum number of turns is reached without invoking finish(), the episode terminates with a max_turns signal.
B.4 Error-Driven Correction
Execution errors (e.g., NameError, SyntaxError) are treated as structured observations rather than fatal failures. The exception message is appended to the context and the model is expected to repair its code in subsequent turns. This enables learning of error recovery behavior during training, consistent with prior work on execution-feedback-driven self-debugging [4].
B.5 System Prompts
The following system prompts are used verbatim during data generation and evaluation.
B.5.1 Base System Prompt
Listing 2: The base CodeAct-style system prompt provided to the agent across all experimental conditions.
You are a CodeAct-style autonomous agent.
You solve tasks by alternating between:
1. Natural-language reasoning (plain text), and
2. Executable simple Python code blob (inside fenced code blocks).
Each step (output) can include at most 1 (one) code block.
Be concise in your reasoning and code.
When you are finished solving the task, ensure that you output a Python
code block which calls the finish tool.
Call the finish tool ONLY after completely solving the task, NOT on every turn.
Execution rules:
- Python code blocks are executed sequentially.
- Only expressions that are printed or explicitly returned are visible to you.
- Variable assignments alone do NOT produce observable output.
- Do not use variable names that conflict with tool names.
Output discipline:
- If a value will be needed for later reasoning or decisions,
you MUST print it (e.g., via print(...)‘) or make it the final expression
in the code block.
- Do not rely on implicit interpreter state visibility.
Tool usage:
- All tool calls must occur inside Python code blocks.
- Do not fabricate tool outputs; rely only on observed execution results.
Error handling:
- If execution fails or a needed value is missing,
explain why and rerun with corrected code.
Completion:
- When the task is complete, provide a final plain-text answer.
- Do not emit further code after completion.
Output Structure:
You must strictly follow this format for every single turn:
1. Reflect upon the previous observation.
2. A single executable Python block.
You prioritize observability and correctness over brevity.
B.5.2 Runtime-State Instructions
To explicitly define the execution contract for the model, the base system prompt is appended with a specific runtime instruction that varies by regime. Listing 3 details the instruction for the persistent environment, while Listing 4 outlines the corresponding reset instruction for the stateless environment.
Listing 3: The runtime instruction appended to the system prompt in the persistent execution regime.
Runtime state: PERSISTENT.
1. Globals persist eternally.
Once you define x = 1‘, it is available forever.
2. NEVER re-import libraries.
3. NEVER paste code from previous steps.
Listing 4: The runtime instruction appended to the system prompt in the stateless execution regime.
Runtime state: RESET.
1. Runtime state resets every turn.
Python variables DO NOT persist.
You must redefine variables and re-import libraries every step.
B.5.3 Few-Shot Demonstrations
In addition to the base system prompt and runtime-state instruction, the agent is provided with a small number of format demonstrations at the beginning of each episode. These demonstrations illustrate the required Reflection–Action–Observation structure and the handling of runtime state. They are intentionally task-agnostic and do not introduce any Opaque Knapsack-specific information; their sole purpose is to make the execution contract concrete. The few-shot content differs between regimes to faithfully reflect the underlying interpreter semantics. Representative demonstrations for each regime are shown below; additional examples follow the same pattern.
Persistent Regime (representative example).
Listing 5: Representative few-shot demonstration for the persistent execution regime. Variables defined in Turn 1 are reused directly in Turn 2 without redefinition.
--- EXAMPLE: PERSISTENT STATE ---
Task: Store items and sum values.
Turn 1
Assistant:
I will initialize the global list `items` and a helper function.
```python
items = [10, 20]
def foo(items):
return items + items
print(len(items))
User: {"observation": {"success": true, "result": null,
[…]
"runtime_state": {"runtime": "persistent",
"active_globals": ["items", "foo"],
"last_step_globals": ["items", "foo"]}}
Turn 2
Assistant:
I can see `items` and `foo` in active_globals, so I reuse them directly.
```python
items = foo(items)
total = sum(items)
print(f"Total: {total}")
finish()
User: {"observation": {"success": true, "result": null,
[…]
"runtime_state": {"runtime": "persistent",
"active_globals": ["items", "foo", "total"],
"last_step_globals": ["items", "foo", "total"]}}
--- EXAMPLE END ---
Stateless Regime (representative example).
Listing 6: Representative few-shot demonstration for the stateless execution regime. State is explicitly printed at the end of Turn 1 and reconstructed from the observation in Turn 2, since active_globals is empty after each reset.
--- EXAMPLE: RESET STATE ---
Task: Store items and sum values.
Turn 1
Assistant:
I will initialize `items` and print its contents so I can retrieve
them next turn, since state will reset.
```python
items = [10, 20]
def foo(items):
return items + items
# State must be printed to survive the reset
print(f"STATE: items={items}")
print(len(items))
User: {"observation": {"success": true, "result": null,
[…]
"runtime_state": {"runtime": "reset",
"active_globals": [],
"last_step_globals": ["items", "foo"]}}
Turn 2
Assistant:
The environment has reset and active_globals is empty. I reconstruct
`items` and `foo` from the previous observation, then compute the total.
```python
# Re-initializing from previous observation
items = [10, 20]
def foo(items):
return items + items
items = foo(items)
total = sum(items)
print(f"Total: {total}")
finish()
User: {"observation": {"success": true, "result": null,
[…]
"runtime_state": {"runtime": "reset",
"active_globals": [],
"last_step_globals": ["items", "foo", "total"]}}
--- EXAMPLE END ---
Role of Few-Shot Examples.
The demonstrations are identical in task complexity and formatting across regimes and do not introduce additional task information. They serve only to illustrate the execution contract enforced by the runtime. The interpreter independently enforces persistence or stateless behavior regardless of the prompt; the demonstrations do not alter the underlying execution semantics. In the mismatch condition, the persistent-trained model is explicitly shown stateless-style demonstrations at evaluation time yet still produces missing-binding errors characteristic of cross-turn state reliance, confirming that the learned dependence is behavioral rather than merely prompted.
B.6 Task-Specific Goal Prompts
For the Opaque Knapsack task, the agent receives the goal and rules detailed in Listing 7.
Listing 7: The task-specific goal prompt appended for the Opaque Knapsack environment.
Goal: Select a subset of items to maximize total value,
subject to a hard capacity constraint.
Rules:
- Do not assume any item properties without inspecting.
- Never take an item unless you have inspected it.
- Never exceed capacity C.
Maintain an explicit running total of current_weight
in a variable and update it immediately after each take.
Appendix C Training Data, Fine-Tuning, and Inference Details
C.1 Training Data Preparation
This appendix describes the preprocessing pipeline that converts raw agent trace logs into chat-formatted fine-tuning examples. The pipeline performs (i) validation, (ii) message extraction, and (iii) context-aware truncation, followed by minimal final sanity checks before writing each example as a JSONL record of the form {"messages": ...}.
The trace list is shuffled once using a fixed RNG seed (we use seed 42); thus, given fixed inputs and settings, the resulting dataset is deterministic. In the experiments summarized in Table 5, we cap the dataset size at 1000 retained examples. Consequently, some traces are unprocessed after the cap is reached; we report both the number of processed traces and the number left unprocessed due to the cap to avoid ambiguity.
All token budgeting and token statistics use the training tokenizer (Qwen3-8B). We set the truncation limit to a conservative budget of L100L-100, where $L$ is the model context length ( $L=16384$ in our runs), to avoid boundary effects during training.
C.1.1 Validation
Each trace is validated before message extraction.
Checks and rejection reasons.
The validator implements the following checks and returns a categorical rejection reason for any failed trace:
  • Minimum normalized score : Reject if score < 0.5.
  • Explicit termination : Reject unless the code of at least one of the last three agent steps contains a call to the "finish()" tool.
  • Repetitive recovery loop detection: Compute textual similarity (difflib.SequenceMatcher ratio) between the most recent assistant message and the previous messages within a sliding window of size 4. Reject if the most recent message is more similar than 0.9 to all earlier messages in the window (i.e., it repeats a near-identical assistant utterance).
  • High bad-error density: For each agent step event, read the reported interpreter execution error. Count an error as bad unless it contains one of the following substrings: "ToolRuntimeException" or "Tool call limit exceeded". Reject if $\texttt{bad errors}/\#\texttt{agent steps}>\texttt{0.1}$.
These outcome- and behavior-driven filters are applied as a fixed configuration across both execution regimes (Table 5). Filtering low-quality or degenerate trajectories reduces the risk that the model imitates behaviors present in low-quality traces [33, 34].
C.1.2 Message Extraction
Validated traces are converted into chat-style message sequences.
From the event stream:
  • All system prompt strings (a list or a single string) are concatenated with blank lines into a single optional system message.
  • The task description is stored as a user message.
  • Each agent step becomes an assistant message (empty/whitespace assistant texts are skipped).
  • For every step except the last, we optionally append a user observation message derived from the interpreter execution result. We serialize a JSON object containing an output field when output is not None, and an error field when error is non-empty. If both fields are absent, no observation message is added for that step.
At load time, we exclude episodes that lack a task or any step events, or that produce fewer than three total messages after extraction.
C.1.3 Context-Aware Truncation
Some extracted trajectories exceed the model’s context limit. We apply structured truncation rather than naive token clipping, since generic head/tail clipping or sliding windows can break causal coherence in long-horizon interactive traces [9]. Let the truncation budget be B=L100B=L-100 tokens where $L$ is the context limit.
Case 0: No truncation needed.
If the full message list fits within $B$:
  • If the final message is an assistant message, retain the sequence unchanged.
  • If the sequence ends with a user observation and the preceding message is assistant, drop the trailing user message so the example ends on an assistant target.
  • Otherwise, discard the trace (fits_but_ends_with_user_unfixable).
Case 1: Truncation required.
If the sequence exceeds $B$, we retain a causally coherent head–tail skeleton and backfill the middle:
  1. 1.
    Head: Always retain the first message, and also retain the second message if it is a user task message. (In the common case with a system prompt, this preserves system + task.)
  2. 2.
    Tail: Enforce that the truncated sequence ends with an assistant message.
    • If the raw sequence ends with assistant, retain the final two messages (subject to not overlapping the head).
    • If the raw sequence ends with user but the preceding message is assistant, retain only that preceding assistant message (dropping the trailing user observation).
    • Otherwise, discard the trace.
  3. 3.
    Feasibility: If the token cost of head + tail alone exceeds $B$, discard the trace.
  4. 4.
    Middle backfill (reverse chronological): Iterate over the remaining middle messages from most recent to oldest, adding each message in full if it fits. If a user message does not fit, attempt to replace its content with the fixed placeholder {"output": "[... Output Omitted for Brevity ...]"} and include it only if the placeholder fits. If the placeholder does not fit, stop backfilling. assistant messages are never masked: if an assistant message does not fit, stop backfilling immediately.
After truncation, we discard any result that is empty or does not end with an assistant message. The preparation procedure also applies a final sanity check that the output contains at least two messages and ends with assistant.
C.1.4 Yield, Skip Reasons, and Token Statistics
Table 5 reports dataset yield and token statistics for the persistent and stateless execution regimes. In both cases, 1,9001{,}900 trace files were available. We cap retention at 1,0001{,}000 examples after shuffling (--max-samples 1000); therefore, some traces are left unprocessed after the cap is reached. “Skipped” counts and skip-reason breakdowns are computed over the processed subset only.
In the processed subsets for these runs, all skipped traces fail validation (rather than message extraction or truncation). Specifically, in the persistent regime, the validator rejects $415$ traces for low score and $15$ for high bad-error density; in the stateless regime, it rejects $250$ for low score, $12$ for zombie loops, and $7$ for high bad-error density.
Table 5: Data preparation configuration and yield statistics for both execution regimes. We discover 1,9001{,}900 traces per regime and shuffle once with a fixed seed (--seed 42). We cap the retained dataset at 1,0001{,}000 examples; “Processed until cap” counts all traces examined before the cap is reached, and “Unprocessed due to cap” counts remaining traces. Token statistics reflect retained, post-truncation examples. Validator settings are identical across runs.
Persistent Stateless
Traces available (found) 1,900 1,900
Processed until cap 1,430 1,269
Retained (training examples) 1,000 1,000
Skipped (within processed) 430 269
Unprocessed due to cap 470 631
Retained / processed (%) 69.93 78.80
Retained / available (%) 52.63 52.63
Min tokens 2,690 5,235
Max tokens 13,008 16,278
Mean tokens 4,630.60 10,761.06
Total tokens 4,630,605 10,761,056
Skipped reasons among processed traces
Score too low 415 250
Repetitive recovery loop 0 12
High error density 15 7
Validator configuration (both runs)
Min trace score 0.5
Max error ratio 0.1
Max loop similarity 0.9
Loop window 4
C.2 Training Details
All models are fine-tuned from the Qwen3-8B [38] base checkpoint using QLoRA [7]. The base model is loaded in 4-bit quantized form (NF4) with double quantization, and computation is performed in bfloat16. We apply Low-Rank Adaptation [15] with rank $r=64$, scaling factor $\alpha=128$, and dropout $0.05$, targeting all attention and MLP projection layers. We train for three epochs using the AdamW [46] optimizer with a cosine learning-rate schedule, peak learning rate $1\times 10^{-4}$, and warmup ratio $0.03$.
The training is done with a maximum sequence length of 16 384 tokens with sample packing enabled. Due to memory constraints at this context length, we use a micro-batch size of 1 with gradient accumulation over 16 steps. Gradient checkpointing and FlashAttention [5] are enabled to reduce memory overhead. All experiments are conducted on a single NVIDIA A100 GPU with fixed random seed.
C.3 Inference Settings
[…]
Appendix D Diagnostic Trace Analysis
This appendix describes the trace-level analysis used to (i) quantify state-mismatch exceptions and (ii) attribute non-optimal outcomes to interpretable failure categories in the cross-evaluation setting.
D.1 State-Mismatch Exception Incidence
We quantify state-mismatch exceptions by scanning each episode for execution errors whose message (lowercased) contains any of the substrings nameerror, unboundlocalerror, or is not defined. These correspond to Python scope errors raised when the agent references a variable absent from the current interpreter state—the expected failure signature when a persistent-trained model is evaluated under a stateless runtime. Only interpreter execution errors are scanned; tool-level and protocol errors are excluded. Multiple matches within a single episode are counted separately.
For a condition with $n$ episodes, we report the total match count $N_{\text{NE}}$, the fraction of episodes containing at least one match, as well as the mean count per episode and per affected episode.
D.2 Termination Taxonomy
Each episode is partitioned by its recorded terminal cause into four categories: budget exhaustion (turn or context limit reached), abnormal termination (unhandled runtime error or tool crash), normal termination (explicit agent-issued finish signal), and other. The specific values assigned to each category are listed in Table 6. Only normally terminated episodes proceed to the failure analysis below, since exhausted or crashed episodes are already attributable to those causes.
Table 6: Terminal cause categories and associated runtime finish signals.
Category Runtime finish signals
Budget exhaustion "max turns", "max steps", "length", "context length"
Abnormal "error", "exception", "tool error"
Normal "finish tool"
Other any other value
D.3 Failure Classification
For normally terminated episodes, we assign exactly one mutually exclusive category using the following decision order:
(1) Optimal.
Episodes with $s=1$ are labeled optimal successes. Applying this check first ensures that episodes containing recoverable execution errors are not misclassified at later stages.
(2) Constraint or protocol violation.
For suboptimal episodes, we apply case-insensitive substring matching against a concatenation of all recorded error messages from the episode. The matched categories are: capacity constraint violations (e.g., exceeds capacity), class constraint violations (e.g., disallowed class), and protocol violations (e.g., must be inspected, already taken). The first matching pattern determines the label.
(3) Execution instability.
If no deterministic violation is detected, we assess error density using two indicators computed over interpreter execution errors only (tool and protocol errors are excluded):
  • Overall error density: $\rho=E/T$, where $E$ is the total number of execution errors and $T$ is episode length in steps, estimated from the trace summary when available and otherwise by counting step events directly (minimum $T=1$).
  • Terminal error concentration: $\tau$, the number of execution errors among the final five recorded events.
An episode is classified as exhibiting execution instability if $\rho>0.5$ or $\tau\geq 3$. This captures trajectories dominated by repeated execution failures, particularly those concentrated late in the episode, where the agent has entered a recovery loop rather than making progress.
(4) Silent suboptimality.
Episodes that terminate normally, match no constraint or protocol violation patterns, do not meet the execution instability thresholds, and achieve $s<1$, are classified as silently suboptimal. These represent feasible but non-optimal solutions reached without any detectable failure signal.
(5) Unclassified.
Any remaining normally terminated episode (e.g., missing score) is labeled unclassified.
Threshold selection.
The instability thresholds ($\rho>0.5$ and $\tau\geq 3$ within the final five events) are fixed heuristic values chosen to separate trajectories dominated by execution errors from cleanly terminating suboptimal runs. A density threshold of 0.5 requires that execution errors constitute the majority of agent steps, reflecting sustained failure rather than isolated mistakes. The terminal concentration criterion ($\tau\geq 3$ of the final five events) captures late-stage error clustering characteristic of recovery loops following state-mismatch exceptions. These thresholds were not tuned to maximize any particular metric or percentage reported in the paper; rather, they were selected a priori to reflect qualitatively distinct failure regimes. In exploratory sensitivity checks, small perturbations of these thresholds did not materially alter the qualitative conclusions.
Appendix E Statistical Significance Tests
This appendix reports the full results of paired two-sided Wilcoxon signed-rank tests used to assess statistical significance of performance differences between execution regimes, as reported in Table 2. All tests are paired by task index, comparing performance on the same task instances across conditions. Tests are conducted separately for Easy and Hard difficulty buckets, each containing 100 tasks. Effect sizes are reported as rank-biserial correlation $r$. Tables 7 and 8 present results for Easy and Hard tasks, respectively.
E.1 Easy Tasks
Table 7 reports all pairwise comparisons for Easy (in-domain) tasks. Both LoRA adapters produce large, highly significant gains over their respective base models ($\Delta>73$ points, $p<10^{-17}$, $r\geq 0.999$).
Table 7: Paired Wilcoxon signed-rank tests for Easy Opaque Knapsack tasks ($n=100$). Means are normalized optimality scores (%). $\Delta$ represents the difference in means (Mean A - Mean B). Both aligned and mismatched fine-tuning yield highly significant performance gains over the base model, but performance differences between the fine-tuned configurations themselves are not statistically significant at this sample size.
Condition A Condition B Mean A Mean B $\Delta$ $p$-value Sig.
Train Semantics Runtime Semantics Train Semantics Runtime Semantics
Trained vs Base
Persistent Persistent vs Base Persistent 81.11 7.73 +73.38 $5.21\times 10^{-18}$ ***
Stateless Stateless vs Base Stateless 81.96 4.01 +77.95 $1.02\times 10^{-17}$ ***
Mismatched Deployment vs Base
Persistent Stateless vs Base Stateless 82.24 4.01 +78.23 $5.84\times 10^{-17}$ ***
Stateless Persistent vs Base Persistent 81.41 7.73 +73.68 $6.15\times 10^{-18}$ ***
Native vs Mismatched Deployment
Persistent Persistent vs Persistent Stateless 81.11 82.24 1.13-1.13 $5.14\times 10^{-1}$ ns
Stateless Stateless vs Stateless Persistent 81.96 81.41 +0.55 $6.18\times 10^{-1}$ ns
End-to-End Persistent vs. End-to-End Stateless
Persistent Persistent vs Stateless Stateless 81.11 81.96 0.85-0.85 $8.21\times 10^{-1}$ ns
Persistent-Trained vs. Stateless-Trained (Fixed Runtime)
Persistent Persistent vs Stateless Persistent 81.11 81.41 0.30-0.30 $9.12\times 10^{-1}$ ns
Persistent Stateless vs Stateless Stateless 82.24 81.96 +0.28 $4.98\times 10^{-1}$ ns
E.2 Hard Tasks
Table 8 reports comparisons for Hard (scaled-difficulty) tasks. The LoRA-vs-base gap remains large and highly significant for both adapters ($\Delta>65$ points, $p<10^{-16}$), confirming that fine-tuning benefits generalize to harder, unseen instances.
Table 8: Paired Wilcoxon signed-rank tests for Hard (scaled-difficulty) Opaque Knapsack tasks ($n=100$). Means are normalized optimality scores (%). $\Delta$ represents the difference in means (Mean A - Mean B). Similar to the Easy split, fine-tuned models significantly outperform the base model, while differences between fine-tuning and deployment runtimes remain directional but not statistically significant.
Condition A Condition B Mean A Mean B $\Delta$ $p$-value Sig.
Train Semantics Runtime Semantics Train Semantics Runtime Semantics
Trained vs. Base
Persistent Persistent vs Base Persistent 75.38 2.24 +73.14 $5.69\times 10^{-18}$ ***
Stateless Stateless vs Base Stateless 67.69 1.79 +65.90 $4.15\times 10^{-17}$ ***
Mismatched Deployment vs Base
Persistent Stateless vs Base Stateless 68.21 1.79 +66.42 $7.82\times 10^{-16}$ ***
Stateless Persistent vs Base Persistent 72.49 2.24 +70.25 $1.22\times 10^{-17}$ ***
Native vs Mismatched Deployment
Persistent Persistent vs Persistent Stateless 75.38 68.21 +7.17 $1.89\times 10^{-1}$ ns
Stateless Stateless vs Stateless Persistent 67.69 72.49 4.80-4.80 $5.62\times 10^{-1}$ ns
End-to-End Persistent vs End-to-End Stateless
Persistent Persistent vs Stateless Stateless 75.38 67.69 +7.68 $1.64\times 10^{-1}$ ns
Persistent-Trained vs. Stateless-Trained (Fixed Runtime)
Persistent Persistent vs Stateless Persistent 75.38 72.49 +2.89 $3.53\times 10^{-1}$ ns
Persistent Stateless vs Stateless Stateless 68.21 67.69 +0.52 $9.27\times 10^{-1}$ ns
Significance levels.
*** $p<0.001$, ** $p<0.01$, * $p<0.05$, ns = not significant.
Appendix F Trace-Level Behavioral Metrics
This appendix reports trace-level metrics that characterize how agents manage and reuse execution state under each training/runtime configuration. These diagnostics complement the outcome-level results in Table 2 and the statistical tests in Appendix E by revealing behavioral differences that are invisible in optimality scores alone (e.g., whether state lives in the interpreter or is redundantly re-expressed in text).
We report metrics separately for Easy (in-domain) and Hard (scaled-difficulty) tasks in Tables 9 and 10.
Computation.
Each episode is a sequence of executed Python blocks (“steps”). Unless stated otherwise, metrics are computed per episode and then averaged over $n=100$ evaluation traces per condition. Total Turns is the sum of executed steps over the 100 traces (i.e., $100\times$ the mean steps/episode shown in Table 2).
Metric definitions.
Metric names are aligned with Figure 3 where applicable.
  • Context Lifespan (y-axis, Figure 3(c)): mean turn-span of user-defined variable names as they appear in the agent’s generated Python across the episode, computed as $(t_{\mathrm{last}}-t_{\mathrm{first}})$ per variable and then averaged. High values indicate that the agent continues to reference the same symbols across many steps in textual code, regardless of whether those bindings remain executable in the interpreter.
  • Imports per Step: mean number of Python import statements (e.g., import x, from x import y) per executed step. Values near 1.0 indicate re-importing on (nearly) every step—a signature of the “amnesia tax” under stateless training.
  • State Utilization (y-axis, Figure 3(a)): mean number of long-range variable reuses per episode, where a reuse is defined as a reference at turn $t^{\prime}\geq t+2$ to a variable defined at turn $t$ with no intervening re-assignment. Non-zero values indicate genuine executable state delegation to the persistent runtime (rather than within-step reuse).
  • Redefinitions per Step: mean number of assignments per executed step to variables that were already present in the interpreter’s active globals at the start of that step. This is non-zero only when the runtime preserves state across turns. High values can reflect either true in-place state updates or entrenched re-derivation habits that persist even when a binding is already available.
F.1 Easy Tasks
[…]
Cross-runtime evaluations reveal which behaviors are learned priors vs. runtime-driven. When the Persistent-trained adapter is run in a Stateless runtime, Context Lifespan remains high (4.69 vs. 2.99 for matched Stateless), indicating continued cross-step symbol reuse in generated code even though the interpreter discards bindings. Imports/Step remains below 1.0 (0.87), suggesting incomplete adaptation to the stateless contract. State Utilization is near-zero (0.01), consistent with the fact that true long-range executable reuse is impossible under a stateless runtime and residual counts are likely coincidental symbol overlap.
Conversely, when the Stateless-trained adapter is run in a Persistent runtime, Interpreter Lifespan becomes non-zero (2.60) and Redefinitions/Step rises (6.78), yet State Utilization remains exactly zero: the agent does not exploit persistent bindings and instead overwrites state each step. Imports/Step stays at 1.00 despite imports being retained, confirming that re-importing is a learned behavior under stateless training rather than a response to runtime necessity.
Table 9: Trace-Level Behavioral Metrics for Easy Opaque Knapsack tasks ($n=100$ traces per condition).
Lifespan (turns)
Train Semantics Runtime Semantics Total Turns Context (text refs) Interpreter (live bindings) Imports per Step State Utilization Redefinitions per Step
Persistent Persistent 350 2.23 2.41 0.29 1.21 7.04
Stateless Stateless 640 2.99 0.00 1.00 0.00 0.00
Persistent Stateless 592 4.69 0.00 0.87 0.01 0.00
Stateless Persistent 525 2.01 2.60 1.00 0.00 6.78
F.2 Hard Tasks
[…]
The persistent-to-stateless mismatch is more extreme on Hard tasks. The Persistent-trained adapter under a Stateless runtime yields the highest total turn count (692) and an elevated Context Lifespan (5.60 vs. 2.40), indicating that the agent continues to carry its persistent-training strategy in symbol usage even when bindings are reset. State Utilization remains near-zero (0.02), again consistent with the impossibility of true executable reuse under stateless execution.
Finally, the Stateless-trained adapter under a Persistent runtime again accumulates live bindings (Interpreter Lifespan = 2.49; Redefinitions/Step = 6.15) without exploiting them (State Utilization = 0.00), and continues to re-import every step (Imports/Step = 1.00), replicating the Easy-task pattern.
Table 10: Trace-Level Behavioral Metrics for Hard Opaque Knapsack tasks ($n=100$ traces per condition).
Lifespan (turns)
Train Semantics Runtime Semantics Total Turns Context (text refs) Interpreter (live bindings) Imports per Step State Utilization Redefinitions per Step
Persistent Persistent 342 2.05 2.32 0.31 1.11 7.01
Stateless Stateless 552 2.40 0.00 1.00 0.00 0.00
Persistent Stateless 692 5.60 0.00 0.86 0.02 0.00
Stateless Persistent 512 1.96 2.49 1.00 0.00 6.15
Appendix G Reproducibility Statement
To ensure the full reproducibility of our empirical findings, we have thoroughly documented our methodology, experimental setup, and data generation pipelines throughout the manuscript and its appendices.
Code Availability: To facilitate full reproducibility, our complete codebase is publicly available at https://github.com/mrcabbage972/agents-learn-runtime. The repository is structured around five key components:
  • Task Instance Generation: The procedural generation pipeline for the Opaque Knapsack environment, used to create the training and evaluation datasets.
  • Teacher Trace Generation: The CodeAct-style agent loop used to produce interleaved reasoning trajectories, including the infrastructure to toggle between persistent and stateless execution semantics.
  • Trace Dataset Analysis: Scripts for calculating trace dataset statistics, including performance and efficiency metrics.
  • Model Training: Axolotl [^1] https://docs.axolotl.ai/ configurations and specific make commands for QLoRA fine-tuning of the Qwen3-8B base model.
  • Inference and Benchmarking: The benchmark harness, vLLM server setup and table generation script, allowing for the reproduction of our main results (Table 2).
Task Generation and Agent Implementation: The complete procedural generation details for the Opaque Knapsack task—including item distributions, budget derivations, and structural rejection sampling constraints—are specified in Appendix A. The exact agent execution loop, tool API, structured runtime-state headers, and verbatim system prompts (including few-shot demonstrations for both regimes) are provided in Appendix B.
Data Preparation and Training: The deterministic preprocessing pipeline used to convert raw teacher traces into chat-formatted fine-tuning examples—including structural validation filters and context-aware truncation logic—is detailed in Appendix C. All fine-tuning and inference hyperparameters are listed in Appendix C.
Evaluation Metrics and Diagnostics: The programmatic heuristics and thresholds used for trace-level failure classification (e.g., unresolved reference errors, execution instability) are formally defined in Appendix D. Full statistical significance test results are reported in Appendix E, and the definitions for all trace-level behavioral metrics (e.g., Interpreter Lifespan, Context Lifespan, State Utilization) are provided in Appendix F.

### v2

@@
-arXiv:2603.01209v1[cs.AI] 01 Mar 2026
+Agents Learn Their Runtime: Interpreter Persistence as Training-Time Semantics
 Abstract
 Tool-augmented LLM agents increasingly solve tasks by interleaving natural-language reasoning with executable Python actions.
-Many agent frameworks equip models with a persistent interpreter, but training traces often leave this execution contract implicit.
-We ask whether interpreter persistence is merely a runtime scaffold, or a training-time semantic that shapes how agents learn to use the interpreter.
-We answer with a controlled $2\times 2$ study, conducted on a single task family and base model, crossing training semantics (fine-tuning on persistent vs.
-stateless-execution traces) and runtime semantics (evaluating in a persistent vs.
+Many agent frameworks equip models with a persistent interpreter, but training traces often leave this runtime assumption implicit.
+We ask whether interpreter persistence is merely a runtime scaffold, or a property of the training data that shapes how agents learn to use the interpreter.
+We answer with a controlled $2\times 2$ study, conducted on a single task family and base model, crossing training conditions (fine-tuning on persistent vs.
+stateless-execution traces) and runtime conditions (evaluating in a persistent vs.
 stateless interpreter).
 We introduce Opaque Knapsack, a deliberately non-collapsible, partially observable knapsack variant with budgeted tool access.
-For each instance, we generate matched trajectories differing only in the persistence/reset execution contract, fine-tune identical models on each variant, and evaluate all four train-runtime combinations.
-We find that misaligning training and runtime semantics produces two characteristic failure modes: a model trained on persistent traces but deployed in a stateless interpreter triggers missing-variable errors in ${\sim}80\%$ of episodes, entering cascading recovery loops that consume its token budget without making progress.
+For each instance, we generate matched trajectories differing only in whether the interpreter state persists or resets between turns, fine-tune identical models on each variant, and evaluate all four train-runtime combinations.
+We find that misaligning training and runtime conditions produces two characteristic failure modes: a model trained on persistent traces but deployed in a stateless interpreter triggers missing-variable errors in ${\sim}80\%$ of episodes, entering cascading recovery loops that consume its token budget without making progress.
 The reverse—a stateless-trained model deployed in a persistent runtime—pays an amnesia tax: it redundantly re-derives state that the interpreter could have retained, using roughly $3.5\times$ more tokens than the aligned persistent condition.
-Notably, these efficiency and stability costs are not accompanied by a statistically significant reduction in solution quality—execution semantics primarily affect how agents reach solutions, not whether they do.
-This suggests that aligning fine-tuning traces with the deployment runtime reduces inference footprint without sacrificing task performance, and that interpreter persistence should be treated as a first-class semantic of agent datasets rather than a hidden implementation detail at runtime.
+Notably, these efficiency and stability costs are not accompanied by a statistically significant reduction in solution quality–interpreter persistence shapes how agents reach solutions, not whether they do.
+This suggests that the runtime used to generate fine-tuning traces should be an explicit design choice, not a hidden implementation detail.
 Keywords Tool-Augmented Language Models $\cdot$ Synthetic Training Data $\cdot$ Training–Inference Alignment
@@
 (2023).
-In these systems, the interpreter is not just a calculator: it is a workspace where the agent can accumulate variables, data structures, and intermediate results across turns.
-This execution contract varies across agent frameworks.
-CodeAct-style agents typically run inside a persistent Python session Wang et al.
-(2024b), while ReAct-style agents externalize state into the context window Yao et al.
-(2023), and other systems maintain dedicated memory stores Packer et al.
-(2024).
-Yet training pipelines usually treat agent traces as token sequences and leave execution semantics implicit.
+In these systems, the interpreter is a workspace where the agent can accumulate variables, data structures, and intermediate results across turns.
+Agent frameworks differ in how they manage interpreter state.
+CodeAct-style agents Wang et al.
+(2024b) run inside a persistent Python session where variables accumulate across turns, but the traces used to post-train the underlying models rarely make this explicit.
 This mismatch motivates our research question:
-Is interpreter persistence merely an inference-time scaffold, or a training-time semantic that shapes how agents learn to use tools?
-The distinction matters in practice: Agent builders routinely fine-tune models on tool-use logs produced by a particular runtime and then deploy the resulting model in a (sometimes different) runtime.
-If execution semantics are learnable, then matching training traces to deployment semantics can determine whether a CodeAct agent exploits persistent state and how efficiently it does so under latency and cost constraints.
-We answer this question with a controlled $2\times 2$ study that disentangles training semantics from runtime scaffolding, conducted on a single task family and base model.
+Is interpreter persistence merely an inference-time scaffold, or a property of training that shapes how agents learn to use tools?
+The distinction matters in practice: models are routinely fine-tuned on traces generated in one runtime and then deployed in another.
+Whether the model absorbed the runtime’s behavior during post-training determines how efficiently and how reliably it operates at deployment.
+We answer this question with a controlled $2\times 2$ study, conducted on a single task family and base model, that disentangles what the model learned during post-training from what the runtime provides at deployment.
 We introduce Opaque Knapsack, a partially observable knapsack variant designed to be non-collapsible for executable-action agents: Item attributes and feasibility constraints are hidden behind budgeted tool calls, and per-step interaction limits force iterative information gathering and state revision.
-For each task instance, we generate paired interleaved reasoning-action trajectories matched by task instance, prompt, tool interface, and supervision, differing only in the persistence/reset execution contract: whether the interpreter state persists across turns or is reset after each action.
+For each task instance, we generate paired interleaved reasoning-action trajectories matched by task instance, prompt, tool interface, and supervision, differing only in whether the interpreter state persists or resets between turns: whether the interpreter state persists across turns or is reset after each action.
 We fine-tune identical Qwen3-8B models on each trace type and evaluate them under both runtime semantics.
@@
 Contributions:
-  • Execution semantics as a training variable.
-We identify interpreter persistence as part of the agent trace contract, studying execution semantics as learnable components of the training signal with a controlled $2\times 2$ train/runtime design.
   • A non-collapsible benchmark with paired trajectories.
@@
   • Evidence that persistence is learned.
-Using a $2\times 2$ train/runtime cross-evaluation, we show that exposure to persistence during fine-tuning shapes state-management behavior.
-These results suggest that interpreter persistence should be treated as a first-class semantic of agent datasets and training protocols, not a hidden implementation detail of the runtime.
-Relatedly, aligning the execution semantics of training traces with the deployment runtime lowers inference cost without compromising performance.
+Using a $2\times 2$ train/runtime cross-evaluation, we show that interpreter persistence is absorbed as a behavioral prior during post-training — carrying into deployment regardless of prompting, and producing characteristic failure modes when training and runtime are misaligned.
 2 Related Work
@@
 However, ACI work largely treats the execution environment as an inference-time systems choice.
-In contrast, we isolate a specific execution semantic—interpreter state persistence—and study it as part of the training-time trace contract, measuring how training/runtime alignment shapes learned state-management behavior.
+In contrast, we isolate a specific execution semantic—interpreter state persistence—and study it as part of the training-time trace semantics, measuring how training/runtime alignment shapes learned state-management behavior.
 Stateful and Non-Collapsible Evaluation Environments.
@@
 Toolformer (Schick et al., 2023) shows that effective tool invocation emerges when tool use is embedded in training, and recent work argues that the allocation and timing of reasoning should align between training and inference (Nguyen et al., 2026; Akter et al., 2025).
-We extend this alignment principle to execution semantics: our results suggest that if persistent state is available at deployment, exposing the same persistence contract during fine-tuning can shape whether (and how) agents learn to delegate state to the interpreter.
+We extend this alignment principle to execution semantics: our results suggest that if persistent state is available at deployment, exposing the same persistence behavior during fine-tuning can shape whether (and how) agents learn to delegate state to the interpreter.
 Conversely, mismatching training trace semantics and deployment runtimes can induce compounding behavioral distribution shift, a known vulnerability in imitation learning (Ross et al., 2011).
@@
 Exact system prompts, runtime headers, and few-shot demonstrations are documented verbatim in Appendix B.
-Few-shot demonstrations and contract consistency.
-Each runtime regime includes a small number of format-only demonstrations whose sole purpose is to make the execution contract concrete (persistent reuse vs.
+Few-shot demonstrations and semantics consistency.
+Each runtime regime includes a small number of format-only demonstrations whose sole purpose is to make the execution semantics concrete (persistent reuse vs.
 mandatory redefinition / explicit printing of state).
@@
 Importantly, the stateless regime is not data-limited (as shown in Table 1): It produces longer, more tool-heavy trajectories and substantially more tokens during trace generation, yet yields a substantially larger interaction footprint and directionally lower solution quality downstream.
-This indicates that token volume is an unreliable proxy for learnability in agent datasets; the execution contract can make traces qualitatively easier or harder to learn from even when tasks, tools, and supervision are held fixed.
-This echoes findings that SFT teaches behavioral style and format as much as task capability (Zhou et al., 2023) — here, the execution contract is the stylistic prior being absorbed.
+This indicates that token volume is an unreliable proxy for learnability in agent datasets; the execution semantics can make traces qualitatively easier or harder to learn from even when tasks, tools, and supervision are held fixed.
+This echoes findings that SFT teaches behavioral style and format as much as task capability (Zhou et al., 2023) — here, the execution semantics is the stylistic prior being absorbed.
 Mismatch failures reveal genuine reliance on executable state.
@@
 These failures are asymmetric: They do not appear in the other train/runtime configurations, indicating that persistent-trained policies are not merely “helped” by statefulness but depend on it as part of their learned execution strategy.
-From an engineering perspective, this suggests a concrete risk: swapping runtimes (or silently changing interpreter persistence) can degrade both correctness and stability without any change to the model weights.
-Implications for dataset and system design.
-Two implications follow: First, execution semantics (e.g., persistence vs.
-statelessness as well as how state is exposed or summarized) should be treated as first-class attributes of agent datasets and training protocols, alongside tool schemata and supervision signals.
-Second, benchmarks for executable-action agents should explicitly prevent collapse into one-shot scripts; Opaque Knapsack provides one such blueprint by combining tool-gated hidden attributes, hard interaction budgets, and non-monotonic constraints that force iterative revision.
-Falsifiability.
-If interpreter persistence were only an inference-time scaffold, we would expect little or no interaction between training semantics and runtime semantics in our design.
-Specifically, a persistent runtime should reduce token cost similarly regardless of which adapter is deployed; a persistent-trained model should transfer cleanly to a stateless interpreter when shown the stateless contract, without systematic missing-binding errors; and the base model should benefit materially from persistent execution once tool access is available.
-Any of these outcomes would weaken the conclusion that persistence is a learned behavioral prior that must be aligned between training traces and the deployment runtime.
-Instead, we observe the opposite in each case: Token efficiency gains are concentrated in the aligned persistent condition, missing-binding errors are asymmetric and specific to the persistent-trained/stateless-runtime mismatch, and the base model remains near-floor regardless of runtime statefulness—which is consistent with persistence functioning as a training-time semantic rather than a post-hoc scaffold.
+This suggests a concrete risk: swapping runtimes (or silently changing interpreter persistence) can degrade both correctness and stability without any change to the model weights.
+Ruling out the scaffold interpretation.
+These results contradict the view that persistence is just a runtime scaffold.
+If it were, a persistent runtime should reduce token cost regardless of how the model was trained, and a persistent-trained model should adapt cleanly when shown stateless demonstrations at inference.
+Neither holds: efficiency gains are concentrated in the aligned persistent condition while missing-binding errors appear only in the persistent-trained/stateless-runtime mismatch.
 Limitations and future work.
@@
 To generate valid trajectories interleaving reasoning and executable actions, we implemented a custom CodeAct-style agent loop.
-This appendix documents the execution contract, runtime alignment across regimes, and the exact system prompts used in the experiments.
+This appendix documents the execution semantics, runtime alignment across regimes, and the exact system prompts used in the experiments.
 B.1 Execution Flow and Turn Structure
@@
 B.5.2 Runtime-State Instructions
-To explicitly define the execution contract for the model, the base system prompt is appended with a specific runtime instruction that varies by regime.
+To explicitly define the execution semantics for the model, the base system prompt is appended with a specific runtime instruction that varies by regime.
 Listing 3 details the instruction for the persistent environment, while Listing 4 outlines the corresponding reset instruction for the stateless environment.
@@
 These demonstrations illustrate the required Reflection–Action–Observation structure and the handling of runtime state.
-They are intentionally task-agnostic and do not introduce any Opaque Knapsack-specific information; their sole purpose is to make the execution contract concrete.
+They are intentionally task-agnostic and do not introduce any Opaque Knapsack-specific information; their sole purpose is to make the execution semantics concrete.
 The few-shot content differs between regimes to faithfully reflect the underlying interpreter semantics.
@@
 The demonstrations are identical in task complexity and formatting across regimes and do not introduce additional task information.
-They serve only to illustrate the execution contract enforced by the runtime.
+They serve only to illustrate the execution semantics enforced by the runtime.
 The interpreter independently enforces persistence or stateless behavior regardless of the prompt; the demonstrations do not alter the underlying execution semantics.